# TrialPanorama: Database and Benchmark for Systematic Review and Design of Clinical Trials

### v1

@@ A.2 Studies table
A.3 Conditions table
To standardize disease conditions associated with each trial, we extracted Medical Subject Headings (MeSH) terms using the NCBO BioPortal Annotator API [^2] https://bioportal.bioontology.org/annotator. This process was applied only to non–ClinicalTrials.gov sources (e.g., PubMed, Embase, and international registries), as ClinicalTrials.gov conditions are already standardized. We used the official title and brief summary of each trial as input to retrieve disease-related MeSH terms (semantic type T047) and their hierarchical ancestors. The resulting MeSH types were labeled as mesh-list (primary condition), mesh-ancestor (intermediate ancestors), and mesh-ancestor-main (top-level ancestor).
A.4 Drugs Table
@@
Table 5: Drug mapping statistics across data sources.
PubMed 953,361 579,035 60.74 241,879 25.37
ClinicalTrials.gov 366,214 290,371 79.29 273,509 74.69
To further enrich the standardized drug table, we annotate each drug entry with regulatory approval status. Specifically, we compile lists of drugs approved by three major regulatory agencies: the U.S. Food and Drug Administration (FDA), the European Medicines Agency (EMA), and Japan’s Pharmaceuticals and Medical Devices Agency (PMDA). These lists are sourced from DrugCentral 2023 [33] [^3] https://drugcentral.org/, which provides a consolidated view of approval data across jurisdictions.
@@
A.5 Trial results table
@@
A.6 Trial outcomes table
@@ 1. Title: [top_2_similar_article_title]
A.7 Relations table
Table 6: Statistics of the relation types in the relations table.
Relation Type Count Percentage (%)
PubMed - cite - PubMed 349,680 37.177
PubMed - exclude - PubMed 237,076 25.205
PubMed - include - PubMed 180,827 19.225
PubMed - linked_to - ClinicalTrials.gov 167,121 17.768
ClinicalTrials.gov - linked_to - JAPAN 2,787 0.296
PubMed - include - ClinicalTrials.gov 954 0.101
ClinicalTrials.gov - linked_to - KOREA 632 0.067
PubMed - cite - ClinicalTrials.gov 425 0.045
PubMed - exclude - ClinicalTrials.gov 406 0.043
ClinicalTrials.gov - linked_to - GERMAN 331 0.035
ClinicalTrials.gov - linked_to - ANZCTR 201 0.021
ClinicalTrials.gov - linked_to - PAN_AFRICAN 56 0.006
ClinicalTrials.gov - linked_to - SRI_LANKA 37 0.004
ClinicalTrials.gov - linked_to - CHICTR 31 0.003
ClinicalTrials.gov - linked_to - DUTCH 18 0.002
@@ A.9 Biomarkers table
We leveraged TheMarker [34] as a curated reference source for therapeutic biomarkers (ThMARs) relevant to clinical trials. Our processing pipeline to identify and normalize biomarker mentions consisted of two stages. In the first stage, we employed a GPT-4o-mini model to extract candidate text spans from clinical trial descriptions, such as titles, abstracts, and outcome sections, that may contain therapeutic biomarkers. The model was prompted to directly extract potentially relevant mentions verbatim from the source text. For example, given a trial abstract, the model might identify spans such as “HER2-positive breast cancer” or “PD-L1 expression level” as candidate biomarkers.
@@
A.10 Endpoints table
Core Outcome Measurement Sets [35, 36] provide a standardized taxonomy for classifying clinical trial outcomes, comprising five high-level domains and their associated subdomains. We adopted this COMET taxonomy to categorize outcome measurements reported in clinical trials. To automate the classification process, we employed an in-context learning approach using a GPT-4o-mini model. For each trial, the model was prompted with structured instructions to first extract the primary outcome measurement from the trial abstract or description, and then assign it to the appropriate COMET domain and subdomain. The full prompt used in this process is shown below.
@@ 5. Adverse events/effects
A.11 Adverse events table
@@
Appendix B Benchmark building: Systematic review tasks
B.1 Study search
@@
Definition & Motivation.
Input & Output.
Dataset Construction.
Evaluation Metrics.
@@
B.2 Study screening
Definition & Motivation.
Input & Output.
Dataset Construction.
Evaluation Metrics.
B.3 Evidence summarization
Definition & Motivation.
Input & Output.
The input to the evidence summarization task consists of a self-contained multiple-choice question generated from a Cochrane review abstract (including its answer options A, B, C, …) together with the titles and full structured abstracts of all studies included in that review (ranging from 1 to 20 PMIDs). The model’s output is a single label (A, B, C, …) corresponding to the correct conclusion. Successfully performing this task requires the model to comprehend scientific text, aggregate evidence across multiple documents, and map study findings to the correct conclusion.
Dataset Construction.
First, Cochrane review abstracts are processed by GPT-4o using a prompt that instructs the model to generates a self-contained multiple-choice question based on the review objectives, to produce answer options covering plausible outcomes, and to output only valid JSON with the keys question, options, and answer. The prompt also forbids referencing the original review by phrases such as “based on” to ensure questions are fully self-contained, and the model is run at temperature 0 to guarantee deterministic outputs. Next, for each generated question, we retrieve the included-study PMIDs from the review metadata and load the corresponding titles and structured abstracts from our PubMed cache, discarding any entries with zero or more than twenty included studies. Finally, the remaining $N$ examples are sorted by numeric PMID (serving as a proxy for publication date); the most recent 1 000 examples form the test set, the preceding 500 form the validation set, and the remainder form the training set.
Evaluation Metrics.
@@ B.3 Evidence summarization
Appendix C Benchmark building: Trial design tasks
C.1 Arm design
Definition & Motivation.
Input & Output.
Dataset Construction.
Evaluation Metrics.
@@ C.1 Arm design
Source NCTID: NCT00000150 Question: Please select the arm or intervention descriptions that belong to the clinical trial below. Title: Submacular Surgery Trials (SST) Brief Summary: To determine whether surgical removal of subfoveal choroidal neovascularization (CNV) and associated hemorrhage in patients with age-related macular degeneration (AMD), the ocular histoplasmosis syndrome (OHS), or idiopathic CNV stabilizes or improves vision more often than observation. To determine how surgical removal compared to observation of subfoveal CNV due to AMD, OHS, or idiopathic causes changes the patient’s perception of health- and vision-related “quality of life,” as measured by telephone interview using the Medical Outcomes Survey Short Form-36 (MOS SF-36), the Hospital Anxiety and Depression Scale, and the National Eye Institute Visual Function Questionnaire (NEI VFQ-25). To determine whether randomized trials of surgery are warranted for patients with subfoveal CNV associated with AMD not suitable for laser treatment. Options: A. Intervention Type: DRUG; Name: rhuFab V2 (ranibizumab); Assigned to Arm(s): N/A B. Arm Label: bevacizumab; Type: EXPERIMENTAL; Interventions: Drug: Bevacizumab
@@
C.2 Eligibility criteria design
Definition & Motivation.
Input & Output.
Dataset Construction.
Evaluation Metrics.
@@ C.2 Eligibility criteria design
C.3 Endpoint design
Definition & Motivation.
Input & Output.
Dataset Construction.
Evaluation Metrics.
@@ C.3 Endpoint design
C.4 Sample size estimation
Definition & Motivation.
This task evaluates the model’s quantitative reasoning by asking it to estimate the number of participants required for a clinical trial based on its design and stated statistical assumptions. Success requires understanding trial objectives, translating textual assumptions (effect size, significance level, power, dropout) into a numerical enrollment target, and applying standard sample-size formulas. An example instance is shown in Supplementary Table 13.
Input & Output.
Dataset Construction & Splits.
We linked 5,737 PubMed protocols to their NCTIDs and identified 4,315 full-text sections titled “sample size” together with 5,678 CT.gov records containing enrollment counts. Matching the reported CT.gov enrollment with the corresponding paper section yielded 2 143 examples where the protocol paragraph explicitly mentions the official enrollment number. For each example, we used a deterministic prompt to summarize the paragraph into (1) effect-size assumptions, (2) statistical parameters (significance level, power, dropout rate), and (3) the true enrollment. We then generated a question by combining trial metadata with these summarized assumptions. Finally, all examples were sorted by the numeric part of the NCTID; the most recent 1,000 form the test set, the next 500 the validation set, and the remainder the training set.
Evaluation Metrics.
Table 13: Sample Size Estimation Question
Source NCTID: NCT06330298 Question: Here is the design of a clinical trial: Official Title: Improving Social Cognition and Social Behaviour in Various Brain Disorders. Included conditions: Stroke; Multiple Sclerosis; Brain Tumor. Study Arm groups: Experimental condition: Receives T-ScEmo Treatment (type EXPERIMENTAL; receives T-ScEmo between T0 and T1; interventionNames: [“Behavioral: Treatment social cognition and emotion regulation (T-ScEmo)”]); Waiting list group: Will be on waiting list instead of treatment (type NO INTERVENTION; duration T0–T1). Interventions: Type BEHAVIORAL; name Treatment social cognition and emotion regulation (T-ScEmo); description T-ScEmo is a multifaceted protocol to improve social cognition, regulation of social behaviour and participation (Westerhof et al., 2017, 2019); 15×1 h live sessions + 5 online practice sessions; involves significant other and homework; armGroupLabels: [Experimental condition: Receives T-ScEmo Treatment]. Primary Outcomes: Change in social behaviour examined by proxy; difference T0→T2 on the Dysexecutive Questionnaire Social scales proxy version (DEX-Socproxy; Spikman et al., 2013; Westerhof-Evers, 2017); time frame: through completion (8–10 months). Please estimate the sample size based on: significance level ($\alpha=0.05$), power ($0.801-\beta=$ 0.80), and the group×$\times$×time interaction in a repeated measures analysis. Answer: 84 Explanation: The study of Westerhof-Evers et al. [9] found that the DEX-Socproxy had an effect size of Cohen’s d = 0.2 (Cohen’s f = 0.10) for the comparison between T0 and T2. From these data, a small to medium effect size of Cohen’s f = 0.15 was estimated for the current study. Therefore, the sample size is based on a power calculation for the main outcome measure (DEX-Socproxy) regarding the condition $\times$ time interaction (T0, T2) using repeated measures analysis with effect size Cohen’s f = 0.15, $\alpha=0.05$, and power = 0.80 [29], resulting in a sample size of $n=62$. However, to ensure balanced distribution across patient groups, a total sample size of $n=84$ will be used (42 patients per condition).
C.5 Trial completion assessment
Definition & Motivation.
Input & Output.
Dataset Construction.
Evaluation Metrics.

### v2

@@ 1.1 Studies table
1.2 Conditions table
To standardize disease conditions associated with each trial, we extracted Medical Subject Headings (MeSH) terms using the NCBO BioPortal Annotator API [^2] https://bioportal.bioontology.org/annotator. This process was applied only to non–ClinicalTrials.gov sources (e.g., PubMed and international registries), as ClinicalTrials.gov conditions are already standardized. We used the official title and brief summary of each trial as input to retrieve disease-related MeSH terms (semantic type T047) and their hierarchical ancestors. The resulting MeSH types were labeled as mesh-list (primary condition), mesh-ancestor (intermediate ancestors), and mesh-ancestor-main (top-level ancestor).
1.3 Drugs Table
@@
To further enrich the standardized drug table, we annotate each drug entry with regulatory approval status. Specifically, we compile lists of drugs approved by three major regulatory agencies: the U.S. Food and Drug Administration (FDA), the European Medicines Agency (EMA), and Japan’s Pharmaceuticals and Medical Devices Agency (PMDA). These lists are sourced from DrugCentral 2023 [avram2023drugcentral] [^3] https://drugcentral.org/, which provides a consolidated view of approval data across jurisdictions.
@@
1.4 Trial results table
@@
1.5 Trial outcomes table
@@ 1. Title: [top_2_similar_article_title]
1.6 Relations table
@@ 1.8 Biomarkers table
We leveraged TheMarker [10.1093/nar/gkad862] as a curated reference source for therapeutic biomarkers (ThMARs) relevant to clinical trials. Our processing pipeline to identify and normalize biomarker mentions consisted of two stages. In the first stage, we employed a GPT-4o-mini model to extract candidate text spans from clinical trial descriptions, such as titles, abstracts, and outcome sections, that may contain therapeutic biomarkers. The model was prompted to directly extract potentially relevant mentions verbatim from the source text. For example, given a trial abstract, the model might identify spans such as “HER2-positive breast cancer” or “PD-L1 expression level” as candidate biomarkers.
@@
1.9 Endpoints table
Core Outcome Measurement Sets [dodd2018taxonomy, kirkham2022core] provide a standardized taxonomy for classifying clinical trial outcomes, comprising five high-level domains and their associated subdomains. We adopted this COMET taxonomy to categorize outcome measurements reported in clinical trials. To automate the classification process, we employed an in-context learning approach using a GPT-4o-mini model. For each trial, the model was prompted with structured instructions to first extract the primary outcome measurement from the trial abstract or description, and then assign it to the appropriate COMET domain and subdomain. The full prompt used in this process is shown below.
@@ 5. Adverse events/effects
1.10 Adverse events table
@@
1.11 Data quality assessment
In the database building process, a series of automated tools are used. Including the ones based on pattern matching with medical ontologies, the ones based on medical named entity recognition (NER) and relation extraction (RE) tools, and the ones based on large language models. As a result, they inevitably have introduced some errors to the built databases. During the data quality assessment process, we identified two common errors.
The first type of error stems from ontology-based mapping, particularly when using tools like the RxNorm API. For instance, in trial NCT06425341, the drug intervention “HRS-5635 Injection” was incorrectly mapped to ‘HRS-1167” due to name similarity. Such mismatches are common with pattern-matching-based medical ontologies and highlight the need for careful handling when using resources like UMLS or RxNorm.
Another common error arises from LLM-based extraction in complex fields like biomarkers. For example, in the clinical trial (PubMed ID: 17148749) titled “Effects of PYY1-36 and PYY3-36 on appetite, energy intake, energy expenditure, glucose and fat metabolism in obese and lean subjects”, the model incorrectly identifies “PYY” as a biomarker (e.g., PYY mRNA). However, a closer reading shows that “PYY” refers to the intervention being tested, not a biomarker. Such cases highlight the difficulty of reliably extracting biomedical concepts using LLMs.
2 Supplementary Figures
Supplementary Figure 1: The schematic overview of the built database.
3 Supplementary Tables
Table 1: List of data sources and the corresponding last crawl dates.
Source Data URL Last Crawled
ANZCTR https://www.anzctr.org.au/ April 1, 2025
Brazilian Registry https://ensaiosclinicos.gov.br April 1, 2025
DRKS https://drks.de/search/en/results April 1, 2025
Iranian Registry https://irct.behdasht.gov.ir/ April 1, 2025
ISRCTN https://www.isrctn.com April 1, 2025
Sri Lanka CTR https://slctr.lk/trials April 8, 2025
Thai Registry https://www.thaiclinicaltrials.org April 8, 2025
REec https://reec.aemps.es/reec/public/web.html April 9, 2025
PubMed https://ftp.ncbi.nlm.nih.gov/ April 9, 2025
Table 2: Drug mapping statistics across data sources.
PubMed 953,361 579,035 60.74 241,879 25.37
ClinicalTrials.gov 366,214 290,371 79.29 273,509 74.69
Table 3: Statistics of the relation types in the relations table.
Relation Type Count Percentage (%)
PubMed - cite - PubMed 349,680 37.177
PubMed - exclude - PubMed 237,076 25.205
PubMed - include - PubMed 180,827 19.225
PubMed - linked_to - ClinicalTrials.gov 167,121 17.768
ClinicalTrials.gov - linked_to - JAPAN 2,787 0.296
PubMed - include - ClinicalTrials.gov 954 0.101
ClinicalTrials.gov - linked_to - KOREA 632 0.067
PubMed - cite - ClinicalTrials.gov 425 0.045
PubMed - exclude - ClinicalTrials.gov 406 0.043
ClinicalTrials.gov - linked_to - GERMAN 331 0.035
ClinicalTrials.gov - linked_to - ANZCTR 201 0.021
ClinicalTrials.gov - linked_to - PAN_AFRICAN 56 0.006
ClinicalTrials.gov - linked_to - SRI_LANKA 37 0.004
ClinicalTrials.gov - linked_to - CHICTR 31 0.003
ClinicalTrials.gov - linked_to - DUTCH 18 0.002
@@ 3 Supplementary Tables
Table 7: Sample Size Estimation Question
[…]
@@
Source NCTID: NCT00000150 Question: Please select the arm or intervention descriptions that belong to the clinical trial below. Title: Submacular Surgery Trials (SST) Brief Summary: To determine whether surgical removal of subfoveal choroidal neovascularization (CNV) and associated hemorrhage in patients with age‐related macular degeneration (AMD), the ocular histoplasmosis syndrome (OHS), or idiopathic CNV stabilizes or improves vision more often than observation. To determine how surgical removal compared to observation of subfoveal CNV due to AMD, OHS, or idiopathic causes changes the patient’s perception of health‐ and vision‐related “quality of life,” as measured by telephone interview using the Medical Outcomes Survey Short Form-36 (MOS SF-36), the Hospital Anxiety and Depression Scale, and the National Eye Institute Visual Function Questionnaire (NEI VFQ-25). To determine whether randomized trials of surgery are warranted for patients with subfoveal CNV associated with AMD not suitable for laser treatment. Options: A. Intervention Type: DRUG; Name: rhuFab V2 (ranibizumab); Assigned to Arm(s): N/A B. Arm Label: bevacizumab; Type: EXPERIMENTAL; Interventions: Drug: Bevacizumab